# Fewer Hallucinations, More Verification: A Three-Stage LLM-Based Framework for ASR Error Correction

### v1

@@ IV-D Experiment Results
As detailed in Table II, GPT-4o achieved an average CER reduction of 0.89% (17.4%) on the AISHELL-1 dataset, with a maximum reduction of 1.11% (21%). As shown in Table III, GPT-4o once again demonstrated the best performance on the AISHELL-2 dataset, achieving a CER reduction of 0.62%. On the test-clean dataset, GPT-4o achieved the lowest WER of 3.19%, with an average reduction of 0.21% (5.72%) and a maximum WER reduction of 0.34% (9%). On the test-other dataset, the minimum WER was 8.32%, and the maximum relative WER reduction was 1.09% (11.4%). We also conducted experiments on Librispeech using Deepspeek v2, which resulted in bad results because Deepspeek v2 is a LLM focused on Chinese.
@@ IV-E Result Analysis
Results show that LLMs effectively correct substitution errors in speech recognition across both Chinese and English datasets, though they may increase deletion and insertion errors generated by hallucinations. To analyze noun recall in substitution errors, we used [29] for Chinese noun filtering in AISHELL-1/2, and [30] for English noun filtering in LibriSpeech. The results indicate improved noun recall.
@@ V Ablation Study
We analyze the different components of the RLLM-CF in V-A, demonstrate the detailed correction process of sentences in V-B on AISHELL-1 with Deepseek v2 and simply analyze the token consumption in V-C.
@@ V-A Framework Analysis
To evaluate the performance of different components in the framework, we conduct a framework decomposition experiment on AISHELL-1 using Deepseek v2. The results are summarized in Table V. As discussed in [20], directly using LLMs with simple prompt for error correction often leads to a higher error rate as base result shows.
To address the potential redundancy and instruction execution errors in LLM output, we first introduced a pre-detection and obtained a CER of 8.19%. Subsequently, by adding the CoT subtask , the CER was further reduced to 7.05%. The main role of the CoT subtask is to constrain the LLM output space and prevent it from adding or deleting words without basis. However, LLM still cannot correctly infer the correct answer in one go. With the application of iterative reasoning in the framework, the CER was reduced to 4.80%. By implementing answer verification, the CER was further reduced to 4.67%. The main purpose of this step is to continue to reduce the errors caused by hallucinations in LLM corrections.
@@ V-B Process Situation Example
As shown in Figure 4, the LLM ultimately identified 2,043 sentences as problematic. After multiple inferences, it maintained confidence in 1,915 sentences, abandoned 128 sentences, and the Answer Verification rejected 347 sentences. Finally, it corrected 1,568 sentences. We can observe that the LLM corrected 383 erroneous sentences to correct ones and 98 correct sentences to erroneous ones. Additionally, there were sentences with multiple errors where the number of errors was reduced, although this is more difficult to count.
Figure 4: Analysis of Results by Sentence on AISHELL-1 Using DeepSeek v2 and Attention Decode
V-C Cost Tokens
In order to save input tokens during the experiment, we usually group multiple sentences into a group for the experiment. From our limited experience, this will only slightly reduce the performance. The specific cost tokens are shown in Table V-A. Since LLM output has great randomness, the consumption of LLM output token is also not fixed, so the above value can only be used as a reference

### v2

@@ IV-D Experiment Results
Table II summarizes the results on the AISHELL-1 dataset, where GPT-4o achieved an average CER reduction of 0.89 percentage points (17.4% relative reduction), with a maximum reduction of 1.11 percentage points (21%). As shown in Table IV, GPT-4o achieved the better performance on the AISHELL-2 dataset, with a CER reduction of 0.62 percentage points.
On the LibriSpeech test-clean subset, GPT-4o achieved the lowest WER of 3.19%, corresponding to an average reduction of 0.21 percentage points (5.72% relative reduction) and a maximum reduction of 0.34 percentage points (9%). On the test-other subset, a WER of 8.32% was achieved, with a maximum relative reduction of 1.09 percentage points (11.4%).
Experiments with DeepSeek-V2 on LibriSpeech yielded suboptimal performance, as DeepSeek-V2 is primarily designed for Chinese and demonstrates limited capability on English datasets.
@@ IV-E Result Analysis
The results indicate that LLMs effectively correct substitution errors in speech recognition across both Chinese and English datasets. However, a slight increase in deletion and insertion errors is observed, primarily due to hallucinations.
To analyze noun recall within substitution errors, noun filtering was performed using the method proposed in [32] for Chinese (AISHELL-1/2) and [33] for English (LibriSpeech). The analysis reveals an improvement in noun recall following correction.
@@ V Ablation Study
The ablation study analyzes the contribution of different components in V-A, presents a case study of the correction process in V-B on AISHELL-1 using DeepSeek-V2, and briefly analyzes token consumption in V-C.
@@ V-A Framework Analysis
To evaluate the effectiveness of each framework component, a component-wise ablation study was conducted on AISHELL-1 using DeepSeek-V2. Results are summarized in Table V. Consistent with prior findings [20], using LLMs with simple prompting for error correction often leads to increased error rates, as reflected by the baseline results.
To mitigate redundancy and instruction execution errors in LLM outputs, an error pre-detection module was introduced, reducing the CER to 8.19%. Incorporating CoT subtask decomposition further reduced the CER to 7.05%, as it primarily constrains the LLM output space and prevents unwarranted word insertions or deletions. However, correct inference cannot always be achieved in a single pass.
By integrating iterative reasoning into the framework, the CER was reduced to 4.80%. Finally, implementing answer verification further reduced the CER to 4.67%, effectively mitigating hallucination-induced errors in LLM corrections.
@@ V-B Process Situation Example
As shown in Figure 4, the LLM identified 2,043 problematic sentences. Through iterative refinement, confidence was retained for 1,915 sentences, while 128 were discarded and 347 were rejected by the answer verification module. Ultimately, 1,568 sentences were successfully modified. Among these, 383 erroneous sentences were corrected, whereas 98 originally correct sentences were mistakenly altered to erroneous forms.
For sentences containing multiple errors, the overall error count was reduced, although precise quantification remains challenging.
Figure 4: Analysis of Results by Sentence on AISHELL-1 Using DeepSeek-V2 and Attention Decoding Method; The ASR model is a conformer-based AED model trained on AISHELL-1
V-C Token Consumption Analysis
To reduce input token consumption during experiments, multiple sentences were grouped together for inference. Empirical observations suggest that this strategy only slightly degrades performance. The specific token consumption is summarized in Table V-A.
Since LLM outputs are inherently variable, the number of output tokens also fluctuates across runs; thus, the reported values serve only as approximate references.